OpenStudy (anonymous):

Which of the following are point-slope equations of the line going through (-4, -7) and (3, 2)? Check all that apply.

OpenStudy (anonymous):

Find slope

OpenStudy (saifoo.khan):

(2+7)/3+4) 9/7

OpenStudy (saifoo.khan):

(2+7)/(3+4) 9/7

OpenStudy (anonymous):

A. y + 7 = -(x + 4) B. y + 2 = -(x + 3) C. y - 2 = (x - 3) D. y - 2 = -(x - 3) E. y + 7 = (x + 4) F. y - 7 = (x - 4)

OpenStudy (saifoo.khan):

are these the options?

OpenStudy (saifoo.khan):

u hav to find the equation right?

OpenStudy (anonymous):

equations.......yes

OpenStudy (saifoo.khan):

OK,then u hav to find the y-intercpet. y = mx + c 2 = 9/7 (3) + c c = -27/7 + 2

OpenStudy (saifoo.khan):

c = -1.85

OpenStudy (anonymous):

A. y + 7 = -(x + 4) B. y + 2 = -(x + 3) C. y - 2 = (x - 3) D. y - 2 = -(x - 3) E. y + 7 = (x + 4) F. y - 7 = (x - 4)